OpenStudy (anonymous):

. A water tank is 60% full. The capacity of the tank is 282 liters. What is the volume of water in the tank?

OpenStudy (anonymous):

You can solve this as a proportion. x/282 = 60/100

OpenStudy (dumbcow):

282*0.6 = answer